Find the moduli and principal arguments of w, z, wz and w/z, given that
(a) w = 10i, z = 1 + i√3, (b) w = −2√3 + 2i, z = 1 − i, (c) w = 1
1 + i√3,
1
1
Expert's answer
2020-05-25T19:05:47-0400

a) |w| = 10, arg(w) = "\frac{\pi}{2}"; |z|= 2, arg(z) = "\frac{\pi}{3}"; wz="10*2*e^{\frac{\pi i}{2}}*e^{\frac{\pi i}{3}} = 20*e^{\frac{5\pi i}{6}} => |wz|=20, arg(wz)=\frac{5\pi}{6}; w/z = 10/2*e^{\frac{\pi i}{6}} => |w/z|=5, arg(w/z)=\frac{\pi}{6}"

b)|w|=4, arg(w)="\frac{2\pi}{3}"; |z| = "\sqrt2", arg(z) = "-\frac{\pi}{4}"; wz = "4*\sqrt2*e^{\frac{5\pi i}{12}} => |wz|=4*\sqrt2, arg(wz) = \frac{5\pi}{12}";

"w/z=4/\sqrt2*e^{\frac{11\pi i}{12}} => |w/z| = 2*\sqrt2, arg(w/z) = \frac{11\pi}{12}"

c)|w| = 2, arg(2) = "\frac{\pi}{3}"; |z| = 1, arg(z) = 0; wz = w*1=w=>|wz|=2, arg(wz)="\frac{\pi}{3}"; w/z=w/1=w=>|w/z|=2, arg(w/z)="\frac{\pi}{3}"